### v1

@@ 4.1 Definitions
$=f((\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r})))+(|V(G)|-1)f(0^{r})=\ell.$
Node $\mathbf{t}$ has children $\mathbf{t_{1},\ldots t_{\delta(t)}}$. By the definition of $\Pi_{Div}$, if $(w^{1},w^{2},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$, there exists $(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1}),\ldots,(w^{1}_{\delta(t)},\ldots w^{r}_{\delta(t)},\ell_{\delta(t)})\in\Pi_{Div}(t_{\delta(t)})$ such that for each $i\in[r]$, $(w^{i},w^{i}_{1},\ldots,w^{i}_{\delta(t)})\allowbreak\in Process_{\mathfrak{C}_{i},\mathcal{D}}$ and $\ell=\ell_{1}$ if $\delta(t)=1$, $\ell={\ell_{1}+\ell_{2}-|V(G)|f(0^{r})}$ if $\delta(t)=2$. Assume that for each $j\in[\delta(t)]$ if $(w^{1}_{j},\ldots,w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$, then for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t_{j}},w^{i}_{j})$-witness $\alpha_{i}^{j}$ such that $Div(S_{\rho_{1}}(\mathcal{D}_{t_{j}},\alpha_{1}^{j}),\allowbreak\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{j}},\alpha_{r}^{j}))=\ell_{j}.$ For each $i\in[r],\ t^{\prime}\in V(\mathcal{D}_{t})$, set
@@
If $\delta(t)=2$, observe that for each $i\in[r]$, $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})\cup S_{\rho_{i}}(\mathcal{D}_{t_{2}},\alpha_{i}^{2})$, where $S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,\allowbreak S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1})\subseteq V(G_{t_{1}})$, $S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2})\subseteq V(G_{t_{2}})$ and ${V(G_{t_{1}})\cap V(G_{t_{2}})=\emptyset}$. So by Lemma 3.2,
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))\\ =Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1}))+Div(S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2}))-|V(G)|f(0^{r})\\ =\ell_{1}+\ell_{2}-|V(G)|f(0^{r})=\ell.\qed$
Observe that if $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$, where $\ell\geq d$ and $w_{i}\in Accept_{\mathfrak{C}_{i},\mathcal{D}}$ for each $i\in[r]$, then, by Lemma 4.9, for each $i\in[r]$, there exists a $(\mathfrak{C}_{i},\mathcal{D})$-witness $\alpha_{i}$ such that $Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))=\ell$, so $(G,S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))\in Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$.
Combining Lemma 4.8 with Lemma 4.9, we get the following corollary.
Corollary 4.10.
This finishes the proof of Theorem 4.6. ∎
@@ 4.3 Dynamic Programming Core for Vertex Cover
We can describe dynamic programming for Vertex Cover from the previous chapter in the terms of dynamic programming core and membership function. For each node $t\in V(\mathcal{D})$, we define $Process_{\mathfrak{C},\mathcal{D}}(t)$ as follows:
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(C_{1}+C_{2},C_{1},C_{2})|C_{1}\in\mathcal{F}_{t_{1}},C_{2}\in\mathcal{F}_{t_{2}},|C_{1}+C_{2}|\leq k\},$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(\rho_{a\rightarrow b}(C^{\prime}),C^{\prime})|C^{\prime}\in\mathcal{F}_{t^{\prime}}\}.$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(C^{\prime},C^{\prime})|C^{\prime}=(c_{1},\ldots,c_{\omega})\in\mathcal{F}_{t^{\prime}},c_{a}=n_{a}\lor c_{b}=n_{b}\},$
@@
Observe that $Process_{\mathfrak{C},\mathcal{D}}(t)$ describes exactly how each entry in $\mathcal{F}_{t}$ is computed from entries in $\mathcal{F}_{t_{1}},\ldots,\mathcal{F}_{t_{\delta(t)}}$. $Process_{\mathfrak{C},\mathcal{D}}(t)$ is empty if and only if $\mathcal{F}_{t}$ is empty, so we can define $Accept_{\mathfrak{C},\mathcal{D}}$ to be $\{0,\ldots,k\}^{\omega}$.
$Accept_{\mathfrak{C},\mathcal{D}}$ can be computed in time $(k+1)^{\omega}$. Moreover, for each $t\in V(\mathcal{D})$, $Process_{\mathfrak{C},\mathcal{D}}(t)$ can be computed in the same time as $\mathcal{F}_{t}$. So $\tau(\mathfrak{C},\mathcal{D})=\mathcal{O}(|V(\mathcal{D})|\cdot(k+1)^{2\omega}\cdot\omega)$.
Observe that for each $t\in V(\mathcal{D})$ for each entry ${(C,C_{1},\ldots C_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)}$ the values of $C_{1},\ldots C_{\delta(t)}$ determine the value of $C$ and $\delta(t)\leq 2$, so $Size_{\mathfrak{C},\mathcal{D}}\leq(k+1)^{2\omega}$.
@@
Let $\alpha$ be a $(\mathfrak{C},\mathcal{D}_{t},w)$-witness and $S\coloneq S_{\rho}(\mathcal{D}_{t},\alpha)$. Let $w_{j}\coloneq\alpha(t_{j})$ and $S_{j}\coloneq S_{\rho}(\mathcal{D}_{t_{j}},\alpha_{j})$ for $j\in[\delta(t)]$. Observe that $S=\bigcup_{j\in[\delta(t)]}S_{j}$ and $\alpha_{j}\coloneq\alpha(t^{\prime})$ for each $t^{\prime}\in V(\mathcal{D}_{t_{j}})$ is a $(\mathfrak{C},\mathcal{D}_{t_{j}},w_{j})$-witnesses. Then by the induction assumption, $S_{j}$ is a vertex cover of $G_{t_{j}}$ and $w_{j}=(c_{1}^{j},\ldots,c_{\omega}^{j})$ where $c_{i}^{j}$ is the number of vertices of color $i$ in $S_{j}$. If $t=t_{1}\oplus t_{2}$ or $t=recolor(t_{1},a\rightarrow b)$, $E(G_{t})=\bigcup_{i\in[\delta(t)]}E(G_{t_{j}})$, so $S$ is a vertex cover of $G_{t}$. Moreover, $S$ uses $\sum_{j\in[\delta(t)]}c_{i}^{j}$ vertices of color $i$ for each $i\in[\omega]$. If $t=addEdges(t_{1},a,b)$ and $(w,w_{1})\in Process_{\mathfrak{C},\mathcal{D}}(t)$, then $c_{a}^{1}=n_{a}$ or $c_{b}^{1}=n_{b}$, so all of $E(G_{t})\setminus E(G_{t_{1}})$ are covered by $S_{1}$. Therefore, $S=S_{1}$ is a vertex cover of $G_{t}$. Moreover, $S$ uses $c_{i}^{1}$ vertices of color $i$ for each $i\in[\omega]$.
Let $S$ be a vertex cover of $G_{t}$, $S_{j}\coloneq S\cap V(G_{t_{j}})$ for each $j\in[\delta(t)]$. Observe that $S_{j}$ is a vertex cover of $G_{t_{j}}$. Then, by the induction assumption, there exist a $(\mathfrak{C},\mathcal{D}_{t_{j}},w_{j})$-witness $\alpha_{j}$ such that $S_{j}=S_{\rho}(\mathcal{D}_{t_{j}},\alpha_{j})$ and $w_{j}=(c_{1}^{j},\ldots,c_{\omega}^{j})$, where $c_{i}^{j}$ is the number of vertices of color $i$ in $S_{j}$. Observe that if there exists a $(\mathfrak{C},\mathcal{D}_{t_{j}},w_{j})$-witness, then $w_{j}\in\mathcal{F}_{t_{j}}$. So, by the definition of $Process_{\mathfrak{C},\mathcal{D}}(t)$ there exists $(w,w_{1},\ldots,w_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)$ such that $w=(c_{1},\ldots,c_{\omega})$, where $c_{i}=\sum_{j\in[\delta(t)]}c_{i}^{j}$ for each $i\in[\omega]$. Let us define for each $t^{\prime}\in V(\mathcal{D}_{t})$,
@@
The direct conclusion of Theorem 4.6 is the following corollary.

### v2

@@ 3.1 Definitions
$=f(\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))+(|V(G)|-1)f(0^{r})=\ell.$
Node $\mathbf{t}$ has children $\mathbf{t_{1},\ldots t_{\delta(t)}}$. By the definition of $\Pi_{Div}$, if $(w^{1},w^{2},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$, there exists $(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1}),\ldots,(w^{1}_{\delta(t)},\ldots w^{r}_{\delta(t)},\ell_{\delta(t)})\in\Pi_{Div}(t_{\delta(t)})$ such that for each $i\in[r]$, $(w^{i},w^{i}_{1},\ldots,w^{i}_{\delta(t)})\allowbreak\in Process_{\mathfrak{C}_{i},\mathcal{D}}$ and $\ell=\ell_{1}$ if $\delta(t)=1$, $\ell={\ell_{1}+\ell_{2}-|V(G)|f(0^{r})}$ if $\delta(t)=2$. Let $t_{j}$ be a child of $t$. By the induction assumption, if $(w^{1}_{j},\ldots,w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$, then for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t_{j}},w^{i}_{j})$-witness $\alpha_{i}^{j}$ such that $Div(S_{\rho_{1}}(\mathcal{D}_{t_{j}},\alpha_{1}^{j}),\allowbreak\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{j}},\alpha_{r}^{j}))=\ell_{j}.$ For each $i\in[r],\ t^{\prime}\in V(\mathcal{D}_{t})$, set
@@
If $\delta(t)=2$, observe that for each $i\in[r]$, $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})\cup S_{\rho_{i}}(\mathcal{D}_{t_{2}},\alpha_{i}^{2})$, where $S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,\allowbreak S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1})\subseteq V(G_{t_{1}})$, $S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2})\subseteq V(G_{t_{2}})$ and $V(G_{t_{1}})\cap V(G_{t_{2}})=\emptyset$}. So by Lemma˜3.8,
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))$
$=Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1}))+Div(S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2}))-|V(G)|f(0^{r})$
$=\ell_{1}+\ell_{2}-|V(G)|f(0^{r})=\ell.\qed$
We would like to briefly explain at this point why the factor $d^{a}$ is stated in the running time of Theorem˜1.4. Integer $a$ stands for the maximum number of children of a node in the given tree decomposition. The reason why this improvement is possible is that during the DP composition, we need to remember for each entry only the best value of diversity and not all $d$ of them, similarly to our proof of Theorem˜1.2. Such a change requires only a single computation for each table entry compared to the $d^{a}$ factor used originally.
4 Example: The Diverse Vertex Cover Problem
In the Vertex Cover problem given a graph $G$ and integer $k$, one asks whether there exists $S\subseteq V(G)$ of size at most $k$ such that $S$ touches all the edges of $G$. Formally, a k-Vertex Cover is a set of pairs $(G,S)$ where $S\subseteq V(G)$ such that $|S|\leq k$ and for each $uv\in E(G)$ $u\in S$ or $v\in S$. In the diverse version of this problem, given a graph $G$ and integers $k,r,d$, one asks whether there exist $r$ vertex covers in $G$, each of size at most $k$, such that their diversity is at least $d$.
We start by showing a dynamic program that, given a cliquewidth decomposition $\mathcal{D}$ of a graph $G$ of width $\omega$, decides whether there exists a k-Vertex Cover of $G$ in time $f(\omega,k)\cdot|V(\mathcal{D})|$.
4.1 Dynamic Programming for Vertex Cover
Let $\mathcal{D}$ be a cliquewidth decomposition of $G$ of width $\omega$. For $t\in V(\mathcal{D})$, let $\mathcal{F}_{t}$ be a set of tuples $(c_{1},\ldots,c_{\omega})$ such that in $G_{t}$ there exist a vertex cover $S$ of size at most $k$ that uses $c_{i}$ vertices of color $i$ for each $i\in[\omega]$. Note that $|\mathcal{F}_{t}|\leq(k+1)^{\omega}$. Our dynamic programming constructs for each node $t$ set $\mathcal{F}_{t}$ as follows:
  1. 1.
    $t=intro(v,a)$
    $\mathcal{F}_{t}=\{C_{\{v\}},C_{\emptyset}\}$
    and $\mathcal{F}_{t}$ can be computed in time $2\omega$.
  2. 2.
    $t=t_{1}\oplus t_{2}$
    Observe that a set $S$ is a vertex cover of $G_{t}$ if and only if it is a union of vertex covers $S_{1}$ of $G_{t_{1}}$ and $S_{2}$ of $G_{t_{2}}$.
    Denote $(x_{1},\ldots,x_{\omega})+(y_{1},\ldots,y_{\omega})=(x_{1}+y_{1},\ldots,x_{\omega}+y_{\omega})$ and $|(x_{1},\ldots,x_{\omega})|=x_{1}+\cdots+x_{\omega}$. Then
    $\mathcal{F}_{t}=\{C_{1}+C_{2}:C_{1}\in\mathcal{F}_{t_{1}},C_{2}\in\mathcal{F}_{t_{2}},|C_{1}+C_{2}|\leq k\}$
    and $\mathcal{F}_{t}$ can be computed in time $(k+1)^{2\omega}\cdot\omega$.
  3. 3.
    $t=recolor(t^{\prime},a\rightarrow b)$
    Then
    $\mathcal{F}_{t}=\{\rho_{a\rightarrow b}(C^{\prime}):C^{\prime}\in\mathcal{F}_{t^{\prime}}\}$
    and $\mathcal{F}_{t}$ can be computed in time $(k+1)^{\omega}\cdot 2$.
  4. 4.
    $t=addEdges(t^{\prime},a,b)$
    Observe that a set $S$ is a vertex cover of $G_{t}$ if and only if it is a vertex cover of $G_{t^{\prime}}$ and contains all vertices of color $a$ or all vertices of color $b$. Let $n_{i}$ be the number of vertices of color $i$ in $G_{t}$. We can compute it by induction for each $i\in[\omega]$ for each $t\in V(\mathcal{D})$ in time $|V(\mathcal{D})|$. Then
    and $\mathcal{F}_{t}$ can be computed in time $(k+1)^{\omega}\cdot 2$.
Theorem 4.1.
$\mathcal{O}(|V(\mathcal{D})|\cdot(k+1)^{2\omega}\cdot\omega).$
Proof.
For each node $t\in V(\mathcal{D})$, set $\mathcal{F}_{t}$ can be computed in time
$max(2\omega,(k+1)^{2\omega}\cdot\omega,(k+1)^{\omega}\cdot 2).$
4.2 From Vertex Cover to Diverse Vertex Cover
We define dynamic programming entries representing the tuples of partial solutions for each node $t$ of the decomposition. Intuitively, these entries should be combinations of entries from $\mathcal{F}_{t}$. For $t\in V(\mathcal{D})$ let $\mathcal{R}_{t}$ be a set of tuples
[…]
Note that $|\mathcal{R}_{t}|\leq(k+1)^{r\omega}$.
  1. 1.
    $t=intro(v,a)$
    $\mathcal{R}_{t}=\{(C^{1},\ldots,C^{r},\ell):\forall_{j\in[r]}\ C^{j}\in\{C_{\{v\}},C_{\emptyset}\},\ell=f(m(C^{1},\ldots,C^{r},v))\},$
    $\mathcal{R}_{t}$ can be computed in time $2^{r}\cdot r\omega$.
  2. 2.
    $t=t_{1}\oplus t_{t}$
    $\mathcal{R}_{t}=\{(C^{1},\ldots,C^{r},\ell):\exists_{(C_{1}^{1},\ldots,C_{1}^{r},\ell_{1})\in\mathcal{R}_{t_{1}},(C_{2}^{1},\ldots,C_{2}^{r},\ell_{2})\in\mathcal{R}_{t_{2}}}\ s.t.\ \forall_{j\in[r]}\ C^{j}=C_{1}^{j}+C_{2}^{j}\land|C^{j}|\leq k,\\ \ell=max\{\ell_{1}+\ell_{2}-|V(G)|f(0^{r}):\exists_{(C_{1}^{1},\ldots,C_{1}^{r},\ell_{1})\in\mathcal{R}_{t_{1}},(C_{2}^{1},\ldots,C_{2}^{r},\ell_{2})\in\mathcal{R}_{t_{2}}}\ s.t.\ \forall_{j\in[r]}C^{j}=C_{1}^{j}+C_{2}^{j}\land|C^{j}|\leq k\}\}.$
    $\mathcal{R}_{t}$ can be computed in time $(k+1)^{2r\omega}\cdot r\omega$.
    Since $V(G_{t_{1}})$ and $V(G_{t_{2}})$ are disjoint, $(C_{1}^{1},\ldots,C_{1}^{r},\ell_{1})$ represents a tuple $S_{1},\ldots S_{r}\subseteq V(G_{t_{1}})$ of diversity $\ell_{1}$ and $(C_{2}^{1},\ldots,C_{2}^{r},\ell_{2})$ represents a tuple $P_{1},\ldots P_{r}\subseteq V(G_{t_{2}})$ of diversity $\ell_{2}$, by Lemma˜3.8 the diversity of $S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r}$ represented by $(C^{1},\ldots,C^{r},\ell)$ is $\ell_{1}+\ell_{2}-|V(G)|\cdot f(0^{r})$. Observe that $(C^{1},\ldots,C^{r})$ may represent multiple tuples of partial solutions in $G_{t}$, but since we are interested in maximizing the diversity, it suffices to keep only the best diversity value among all these tuples.
  3. 3.
    $t=recolor(t^{\prime},a\rightarrow b)$
    $\mathcal{R}_{t}=\{(\rho_{a\rightarrow b}(C^{1}),\ldots,\rho_{a\rightarrow b}(C^{r}),\ell):(C^{1},\ldots,C^{r},\ell)\in\mathcal{R}_{t^{\prime}}\}.$
    $\mathcal{R}_{t}$ can be computed in time $(k+1)^{r\omega}\cdot 2r$.
  4. 4.
    $t=addEdges(t^{\prime},a,b)$
    $\mathcal{R}_{t}=\{((c_{1}^{1},\ldots,c_{\omega}^{1}),\ldots,(c_{1}^{r},\ldots,c_{\omega}^{r}),\ell)\in\mathcal{R}_{t^{\prime}}:\forall{j\in[r]}\ c_{a}^{j}=n_{a}\lor c_{b}^{j}=n_{b}\}.$
    $\mathcal{R}_{t}$ can be computed in time $(k+1)^{r\omega}\cdot 2r$.
Theorem 4.2.
$\mathcal{O}(|V(\mathcal{D})|\cdot(k+1)^{2r\omega}\cdot r\omega).$
Proof.
For each $\in V(\mathcal{D}), t\mathcal{R}_{t}$ can be computed in time
@@ 4.3 Dynamic Programming Core for Vertex Cover
We describe dynamic programming for Vertex Cover from Section˜4.1 in terms of the dynamic programming core and membership function. For each node $t\in V(\mathcal{D})$, we define $Process_{\mathfrak{C},\mathcal{D}}(t)$ as follows:
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(C_{1}+C_{2},C_{1},C_{2}):C_{1}\in\mathcal{F}_{t_{1}},C_{2}\in\mathcal{F}_{t_{2}},|C_{1}+C_{2}|\leq k\},$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(\rho_{a\rightarrow b}(C^{\prime}),C^{\prime}):C^{\prime}\in\mathcal{F}_{t^{\prime}}\}.$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(C^{\prime},C^{\prime}):C^{\prime}=(c_{1},\ldots,c_{\omega})\in\mathcal{F}_{t^{\prime}},c_{a}=n_{a}\lor c_{b}=n_{b}\},$
@@
Observe that $Process_{\mathfrak{C},\mathcal{D}}(t)$ describes exactly how each entry in $\mathcal{F}_{t}$ is computed from entries in $\mathcal{F}_{t_{1}},\ldots,\mathcal{F}_{t_{\delta(t)}}$. $Process_{\mathfrak{C},\mathcal{D}}(t)$ is empty if and only if $\mathcal{F}_{t}$ is empty, so we can define $Accept_{\mathfrak{C},\mathcal{D}}$ to be $\{0,\ldots,k\}^{\omega}$. $Accept_{\mathfrak{C},\mathcal{D}}$ can be computed in time $(k+1)^{\omega}$. Moreover, for each $t\in V(\mathcal{D})$, $Process_{\mathfrak{C},\mathcal{D}}(t)$ can be computed in the same time as $\mathcal{F}_{t}$. So $\tau(\mathfrak{C},\mathcal{D})=\mathcal{O}(|V(\mathcal{D})|\cdot(k+1)^{2\omega}\cdot\omega)$. Observe that for each $t\in V(\mathcal{D})$, for each entry $(C,C_{1},\ldots C_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)$}, the values of $C_{1},\ldots C_{\delta(t)}$ determine the value of $C$ and $\delta(t)\leq 2$, so $Size_{\mathfrak{C},\mathcal{D}}\leq(k+1)^{2\omega}$.
@@
Let $\alpha$ be a $(\mathfrak{C},\mathcal{D}_{t},w)$-witness and $S\coloneq S_{\rho}(\mathcal{D}_{t},\alpha)$. We will prove that $w=(c_{1},\ldots,c_{\omega})$ and $S$ is a vertex cover of $G$ that uses $c_{i}$ vertices of color $i$. Let $w_{j}\coloneq\alpha(t_{j})$ and $S_{j}\coloneq S_{\rho}(\mathcal{D}_{t_{j}},\alpha_{j})$ for $j\in[\delta(t)]$. Observe that $S=\bigcup_{j\in[\delta(t)]}S_{j}$ and $\alpha_{j}\coloneq\alpha|_{V(\mathcal{D}_{t_{j}})}$ is a $(\mathfrak{C},\mathcal{D}_{t_{j}},w_{j})$-witnesses. Then by the induction assumption, $S_{j}$ is a vertex cover of $G_{t_{j}}$ and $w_{j}=(c_{1}^{j},\ldots,c_{\omega}^{j})$ where $c_{i}^{j}$ is the number of vertices of color $i$ in $S_{j}$. If $t=t_{1}\oplus t_{2}$ or $t=recolor(t_{1},a\rightarrow b)$, $E(G_{t})=\bigcup_{j\in[\delta(t)]}E(G_{t_{j}})$, so $S$ is a vertex cover of $G_{t}$. Moreover, $S$ uses $\sum_{j\in[\delta(t)]}c_{i}^{j}=c_{i}$ vertices of color $i$ for each $i\in[\omega]$. If $t=addEdges(t_{1},a,b)$ and $(w,w_{1})\in Process_{\mathfrak{C},\mathcal{D}}(t)$, then $c_{a}^{1}=n_{a}$ or $c_{b}^{1}=n_{b}$, so all of $E(G_{t})\setminus E(G_{t_{1}})$ are covered by $S_{1}$. Therefore, $S=S_{1}$ is a vertex cover of $G_{t}$. Moreover, $S$ uses $c_{i}^{1}=c_{i}$ vertices of color $i$ for each $i\in[\omega]$.
Now let $S$ be a vertex cover of $G_{t}$ that uses $c_{i}$ vertices of color $i$ for each $i\in[\omega]$. We will prove that there exists a $(\mathfrak{C},\mathcal{D}_{t},w)$-witness $\alpha$ for $w=(c_{1},\ldots,c_{\omega})$ such that $S=S_{\rho}(\mathcal{D}_{t},\alpha)$. Let $S_{j}\coloneq S\cap V(G_{t_{j}})$ for each $j\in[\delta(t)]$. Observe that $S_{j}$ is a vertex cover of $G_{t_{j}}$. Then, by the induction assumption, there exist a $(\mathfrak{C},\mathcal{D}_{t_{j}},w_{j})$-witness $\alpha_{j}$ such that $S_{j}=S_{\rho}(\mathcal{D}_{t_{j}},\alpha_{j})$ and $w_{j}=(c_{1}^{j},\ldots,c_{\omega}^{j})$, where $c_{i}^{j}$ is the number of vertices of color $i$ in $S_{j}$. Observe that if there exists a $(\mathfrak{C},\mathcal{D}_{t_{j}},w_{j})$-witness, then $w_{j}\in\mathcal{F}_{t_{j}}$. So, by the definition of $Process_{\mathfrak{C},\mathcal{D}}(t)$ there exists $(w,w_{1},\ldots,w_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)$ such that $w=(c_{1},\ldots,c_{\omega})$, where $c_{i}=\sum_{j\in[\delta(t)]}c_{i}^{j}$ for each $i\in[\omega]$. Let us define for each $t^{\prime}\in V(\mathcal{D}_{t})$,
@@
The direct conclusion of Theorem˜3.6 is the following corollary.